OpenStudy (leahmarieed):

The side of a square is Two raised to the seven halves power inches. Using the area formula A = s2, determine the area of the square.

OpenStudy (leahmarieed):

A = 196 square inches A = 128 square inches A = 49 square inches A = 4 square inches

OpenStudy (leahmarieed):

will medal and fan

OpenStudy (p0sitr0n):

here, s=2^(7/2) So, A=s^2=(2^(7/2))^2=2^7 inches^2

OpenStudy (p0sitr0n):

or 128 inches^2

OpenStudy (leahmarieed):

thank you so much